OpenStudy (dls):

Show that the ratio of sum of first n terms of GP to the sum of terms from (n+1)th term to (2n)th term is 1/r^n

Parth (parthkohli):

\[S_n = a_1 \left(\dfrac{1 - r^n}{1 - r}\right)\]

Parth (parthkohli):

\[a_{n + 1} = a_1 \cdot r\]

OpenStudy (dls):

\[\LARGE S_n=\frac{a(r^n-1)}{r-1}\] but what about 2nd one

OpenStudy (dls):

how many terms in 2nd GP?

Parth (parthkohli):

\[S= a_1 r\dfrac{}{}\]From \(n + 1\) to \(2n\), there are always \(n\) numbers.

OpenStudy (dls):

why?

Parth (parthkohli):

I mean\[S = a_1 r \left(\dfrac{1 - r^n}{1 - r}\right)\]And that "why?" has an easy answer. Dikhata hoon :-)

Parth (parthkohli):

From \(1\) to \(n\), there are \(n\) numbers.